Vappur density of a compound is39 it contains carbon and hydrogen atoms inthe ratio 1:1 its molecular formula is
 

Grade:11

1 Answers

Nandini kapoor
6 years ago
Molar mass=2×vapour densityThus molar mass of compound is 2×39=78 gramsCH will be the empirical formula(1|1 is the given ratio)Therefore ,equal no. of atoms of C and H are presentEmpirical formula mass is 12+1=13gms.n=molar mass÷empirical formula massso,n=78÷13=6Molecular formula is n×empirical formula=6×CH=C6H6.